OpenStudy (anonymous):

Is there a difference between: ((1 + h)^2 -1^2) / ((1 + h) -1) and ((1 + h)^2 -1) / ((1 + h) -1) Is seems to me that 1^2 should just be simplified to 1. Am I wrong?

OpenStudy (anonymous):

No you are not , In the first case the power here is applied to all the bracket (1+h)^2 = (1^2+h^2+2h) While in the second case ,The power is only applied on the 1 , -1^2= -1 Note: If it was (-1)^2=1

OpenStudy (anonymous):

Good! Thanks Apparently ((1 + h)^2 -1) / ((1 + h) -1) is supposed to be h + 2 Do you know how I get there?